OpenStudy (anonymous):

Rewrite ^3 sqrt 27x-81-5 to make it easy to graph a translation. Describe the graph.

OpenStudy (anonymous):

Please help!

OpenStudy (campbell_st):

well factor the terms inside the cube root \[\sqrt[3]{27(x - 3)} - 5\] what do you know about 27... its 3 to what power...?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

wait no, 3

OpenStudy (campbell_st):

great so the question is now \[\sqrt[3]{3^3(x - 3)} - 5\] can you now simplify the cube root..?

OpenStudy (anonymous):

27x-86

OpenStudy (anonymous):

@campbell_st ^

OpenStudy (campbell_st):

well it can be written as \[3\sqrt[3]{x - 3} -5\]

OpenStudy (anonymous):

so that's the final answer?

OpenStudy (campbell_st):

that would be my final answer